OpenStudy (anonymous):

f(x)=e^(1/x)...i need help explaining the graph and how the exponent affects the graph's different portions so when x<0, 00.

OpenStudy (anonymous):

did you graph it?

OpenStudy (anonymous):

for big values of x \[\frac{1}{x}\] is small (closet to zero) \[e^{\frac{1}{x}}\]gets close to \[e^0=1\]

OpenStudy (anonymous):

likewise as \[x\rightarrow -\infty\] \[\frac{1}{x}\rightarrow 0\] and so again you get 1, meaning you have a horizontal asymptote at y = 1

OpenStudy (anonymous):

not defined at 0, but for small (close to zero) values of x, \[\frac{1}{x}\] gets large and therefore \[e^{\frac{1}{x}}\] gets really large

OpenStudy (anonymous):

you can see the nice picture here http://www.wolframalpha.com/input/?i=y%3De^%281%2Fx%29

OpenStudy (anonymous):

i guess i should add that if x ->0 from the left you get large large negative numbers, so \[lim_{x\rightarrow0^-}e^{\frac{1}{x}}=0\]